OpenStudy (anonymous):

a mp3 player costs a store 270. if the markup is 23 percent what is the selling price

OpenStudy (aaronq):

the price is itself plus 23% price + 0.23*price = price(1+0.23)=1.23*price= (1.23)*270=$ 332.1